# DESI Early Data Release Milky Way Survey Value-Added Catalogue

### v1

@@ Abstract
keyword1 – keyword2 – keyword3
@@ 1 Introduction
In this contribution, we describe the first catalogue of measurements from the stellar part of the Dark Energy Spectroscopic Instrument (DESI) Survey’s Early Data Release (EDR; DESI Collaboration et al., 2023b). DESI is a 5020-fiber, $3.2^{\circ}$-diameter field-of-view, moderate-resolution spectrograph ($50002000<\lambda/\Delta\lambda<$ 5000 for 3600Å $<\lambda<$ 9800Å  spread across three arms) mounted at the prime focus of the Mayall 4m telescope at the Kitt Peak National Observatory (DESI Collaboration et al., 2016; Abareshi et al., 2022). While DESI’s primary mission is to provide sub-percent cosmological measurements of the expansion history of the universe, the instrument is also being used for a Milky Way Survey (DESI-MWS; Cooper et al., 2023) of $\approx 7$ million stars with $r\lesssim 19$ mag, focused on the thick disc and stellar halo of our Galaxy. DESI-MWS will also create the largest spectroscopic catalogue to date of white dwarfs, stars within 100 pc of the Sun, RR Lyrae and BHB stars up to distances of 100 kpc. Serendipitous discoveries have already been made using DESI-MWS data, with the identification of metal-poor stars and rare magnetic white dwarfs exhibiting emission features (Allende Prieto et al., 2023; Manser et al., 2023)
Following the first light with the instrument in October 2020 and the initial period of commissioning, DESI moved into a Survey Validation (SV) phase (DESI Collaboration et al., 2023a). During this period (approximately Dec 2020 to June 2021), the DESI Collaboration observed specific areas of sky (referred to as tiles) with a variety of instrument configurations and target selection schemes to validate observing and data reduction procedures and to optimise the definition of the primary survey target classes (i.e., Milky Way Stars and white dwarfs, Allende Prieto et al. 2020; Cooper et al. 2023; Bright Galaxies, Ruiz-Macias et al. 2020; Hahn et al. 2022; Luminous Red Galaxies, Zhou et al. 2020, 2023; Emission-Line Galaxies, Raichoor et al. 2020, 2023; and QSOs, Yèche et al. 2020; Chaussidon et al. 2023). The final definitions of all the DESI target classes are summarised in Myers et al. (2023). The SV phase was split into 3 campaigns: the first two took an iterative approach to optimise the target selection and survey strategy. During the third campaign, DESI conducted a “One-Percent Survey”, covering approximately 140 deg2 of sky (i.e., 1 per cent of the overall DESI footprint) using the same target selection as the main DESI survey but with much higher completeness across different target classes. After completing the SV phase, DESI began its main 5-year survey program in May 2021.
@@ 2.1 Targets
Details of the DESI target selection and different target classes are given in Myers et al. (2023); DESI Collaboration et al. (2023a, b); Cooper et al. (2023). Here, we provide a summary relevant to Milky Way science and necessary for the interpretation of the value-added catalogue.
@@ 2.2 High-level DESI target types
The DESI_TARGET bitmask encodes the overall DESI survey component for which each source was observed and also separates science targets from flux calibration standard stars and sky fibres. Table 3 lists the subset of DESI_TARGET bits corresponding to standard stars, MWS targets and secondary targets. These bits have the same meaning for SV1, SV2 and SV3, although the actual selections did change somewhat [^2] The magnitude range for STD_BRIGHT was changed from $1815\leq G<$ 18 in SV1 to $1816\leq G<$ 18 in SV2 and SV3. The colour criteria and the magnitude range of STD_FAINT were the same in SV1, SV2 and SV3.. The DESI_TARGET bitmask for a given SV stage is nonzero only for stars observed in that stage. For example, only stars observed in SV1 have a nonzero value of SV1_DESI_TARGET [^3] The VAC also has a small number (759) of rows with zero SV1_DESI_TARGET, SV2_DESI_TARGET and SV3_DESI_TARGET, but non-zero DESI_TARGET. These are fibres targeted as sky but having significant flux in them..
@@
STD_FAINT 33 Standard stars for dark/gray conditions
@@
GAIA_STD_FAINT 33 Gaia-based standard stars for dark/gray conditions
@@ 2.3 MWS survey targets
  • MWS_MAIN_BROAD class correspond to a magnitude limited selection in the range $1916<r<$ 19 without kinematic or colour cuts, supplemented by basic photometric quality requirements and was used in SV1 only  [^4] Several target classes (e.g. MWS_BROAD) have NORTH and SOUTH sub-categories (e.g. MWS_BROAD_NORTH and MWS_BROAD_SOUTH). Sources selected using Bock/Mosaic photometry in the northern section of the DESI survey footprint have the NORTH bit set, while those selected using DECam photometry in the southern section of the footprint have the SOUTH bit set. Since the MWS selection criteria were identical for the northern and southern parts of the footprint, these sub-classes only indicate the origin of the input photometry..
@@
  • MWS_BHB comprises blue horizontal branch stars selected using the DESI Legacy Survey photometry and was used from SV1. We note that this category, which is described in (Cooper et al., 2023), is distinct from the BHB secondary targets category defined for observations of more distant BHB candidates in the dark time (see below).
@@
For all of the target classes listed above in Tables  6, 7, 8 we provide the number of objects in these classes observed in different surveys and programmes included in this VAC. Note that the same object may have more than one target class and may have been observed in multiple surveys.
In this paper we do not describe the selection function of the survey in detail, partially because this VAC is based on the early data release which is a combination of commissioning and science verification surveys and the target selection was changing throughout. However for each survey and programme the information needed to reproduce the selection function or replay the target selection is available as part of the DESI EDR. Specifically:
@@ 2.4 Secondary targets
  • MWS_CALIB and BACKUP_CALIB are classes used during SV only to prioritise stars with data from one or more previous surveys, including APOGEE (Abdurro’uf et al., 2022), BOSS (Dawson et al., 2013), GALAH (De Silva et al., 2015) and Gaia-ESO (Gilmore et al., 2022). The MWS_CALIB stars have Gaia $G$-band mangitudes $1916<G<$ 19 while the BACKUP_CALIB stars have $G<16$.
@@ 2.5 Dedicated fields
Blank 98.0063 -4.8356  [^7] The field was designed to point at the Rosette Nebula, but was mistakenly centered at a negative declination of $\delta=-4.8$ instead of $\delta=4.8$. 1800.3 80718
@@ 4.1 RVS data products
We start by describing the individual RVS outputs of fitting results on coadded DESI spectra. The output of RVS (for a given survey, program and HEALPix pixel) consists of two files: an rvtab file that tabulates measurements by RVSpecFit and an rvmod file that contains the best-fit model for each spectrum.
@@
  • RVS_WARN is an integer bitmask set to zero if there are no warnings during the RVS fit. The first bit (or bit number zero) of the bitmask is set if the CHISQ_TOT - CHISQ_C_TOT $<$ 50, which means that the stellar fit provides only a small improvement over the continuum-only fit. The second bit of RVS_WARN is set to one if the best-fit velocity is within 5 $\mathrm{\,km\,s^{-1}}$ of the radial velocity grid edges ($\pm$1500 $\mathrm{\,km\,s^{-1}}$); the third bit is set to one if the radial velocity error value is larger than 100 $\mathrm{\,km\,s^{-1}}$. The remaining bits of the bitmask are unused in this VAC.
@@ 5.3 Stellar parameter distribution
In Figure 5, we show the distributions of surface gravity and effective temperature (Kiel diagrams) for stars in the VAC, as measured by the RVS (left) and SP pipelines (right). The distributions are similar and show features that are expected, such as the red giant branch (extending from $\log\,g=4$, $T_{\mathrm{eff}}=5500$ K to $\log\,g=1$, $T_{\mathrm{eff}}=4000$ K, the blue horizontal branch ($\log\,g$ $\sim$3, $T_{\mathrm{eff}}>6,500$ K) and the main sequence ($\log\,g$ <4, $T_{\mathrm{eff}}$ <6,000 ). The shape of the main sequence below the turn-off, however, is not very realistic and is likely dominated by systematic effects from the stellar atmospheric grids used in our pipelines. Both pipelines show clustering around grid points in $\log\,g$ and $T_{\mathrm{eff}}$. In addition, the SP pipeline also shows problematic clustering at $T_{\mathrm{eff}}=5000$ K and $\log\,g=5$. Some of these SP artefacts mostly disappear when using stricter selections described in Appendix B.
@@
Figure 6: The distribution of iron abundances for stars in the the MWS VAC as determined by the RVS and SP pipelines. The distribution shows the peak associated with thick disk stars at $\mathrm{\,[Fe/H]}$ =-0.5, the peak associated with GSE at $\mathrm{\,[Fe/H]}=-1.2$ and metal-poor stellar halo at lower metallicities. The distributions are shown with the bin size of 0.1 dex. Due to gridding issue associated with the RVS measurements, would the plot be made with much smaller bins (i.e. 0.01 dex), we would see peaks at location of template grid with the step of 0.25 dex.
@@ 5.6 Draco dwarf galaxy dedicated fields
By looking at the metallicity distribution of Draco member stars in DESI, we find a median metallicity of $\mathrm{\,[Fe/H]}\approx-2.25\pm 0.03$, which at face value seems to differ significantly (by $\sim 0.3$  dex) from the median metallicity of $-1.93\pm 0.01$ reported by Kirby et al. (2011). However Kirby et al. (2011) measured a significant metallicity gradient in Draco of $\approx-0.73$ dex deg-1. The median angular distance of the member stars in the DESI sample from the centre of Draco is $\sim$ 0.25 degrees, as opposed to $\sim$ 0.1 degrees for the literature sample, owing to the large field of view of DESI. This metallicity gradient can, therefore, explain at least $\sim 0.1$  dex of the discrepancy between Kirby et al. (2011) and the DESI measurement. In fact, the median metallicity of the DESI members within 0.1 degrees is $-1.97$  dex perfectly consistent with previous measurements.
@@ Acknowledgments
All the data shown in the published graphs or necessary to reproduce them are available in a machine-readable form on Zenodo https://zenodo.org/10.5281/zenodo.10498911. The value-added catalogue itself is available at https://data.desi.lbl.gov/doc/releases/edr/vac/mws/.
@@ Appendix A Selection criteria of the main MWS targets in SV3
In this appendix we provide the selection criteria for the SV3 targets. The selection code is also available at https://desitarget.readthedocs.io/en/latest/_modules/desitarget/sv3/sv3_cuts.html. We remark that the EDR covers DESI observations prior to the main survey. These targets were selected using Gaia DR2 photometry and astrometry, while the main survey selection uses Gaia DR3 astrometry. Also the selection criteria is slightly different from the main survey selection criteria described in Cooper et al. (2023).
@@
SUCCESS Bit indicating whether the code has likely produced useful results
@@
HEALPIX HEALPixel containing this location at NSIDE=64 in the NESTED scheme

### v2

@@ Abstract
catalogues – surveys – Galaxy: general – Galaxy: kinematics and dynamics – stars: abundances
@@ 1 Introduction
In this contribution, we describe the first catalogue of measurements from the stellar part of the Dark Energy Spectroscopic Instrument (DESI) Survey’s Early Data Release (EDR; DESI Collaboration et al., 2024b). DESI is a 5020-fiber, $3.2^{\circ}$-diameter field-of-view, moderate-resolution spectrograph ($50002000<\lambda/\Delta\lambda<$ 5000 for 3600Å $<\lambda<$ 9800Å spread across three arms) mounted at the prime focus of the Mayall 4 m telescope at the Kitt Peak National Observatory (DESI Collaboration et al., 2016; Abareshi et al., 2022). While DESI’s primary mission is to provide sub-percent cosmological measurements of the expansion history of the universe, the instrument is also being used for a Milky Way Survey (DESI-MWS; Cooper et al., 2023) of $\approx 7$ million stars with $r\lesssim 19$ mag, focused on the thick disc and stellar halo of our Galaxy. DESI-MWS will also create the largest spectroscopic catalogue to date of white dwarfs, stars within 100 pc of the Sun, RR Lyrae and BHB stars up to distances of 100 kpc. Serendipitous discoveries have already been made using DESI-MWS data, with the identification of metal-poor stars and rare magnetic white dwarfs exhibiting emission features (Allende Prieto et al., 2023; Manser et al., 2023)
Following the first light with the instrument in October 2020 and the initial period of commissioning, DESI moved into a Survey Validation (SV) phase (DESI Collaboration et al., 2024a). During this period (approximately Dec 2020 to June 2021), the DESI Collaboration observed specific areas of sky (referred to as tiles) with a variety of instrument configurations and target selection schemes to validate observing and data reduction procedures and to optimise the definition of the primary survey target classes (i.e., Milky Way Stars and white dwarfs, Allende Prieto et al. 2020; Cooper et al. 2023; Bright Galaxies, Ruiz-Macias et al. 2020; Hahn et al. 2023; Luminous Red Galaxies, Zhou et al. 2020, 2023; Emission-Line Galaxies, Raichoor et al. 2020, 2023; and QSOs, Yèche et al. 2020; Chaussidon et al. 2023). The final definitions of all the DESI target classes are summarised in Myers et al. (2023). The SV phase was split into 3 campaigns: the first two took an iterative approach to optimise the target selection and survey strategy. During the third campaign, DESI conducted a “One-Percent Survey”, covering approximately 140 deg2 of sky (i.e., 1 per cent of the overall DESI footprint) using the same target selection as the main DESI survey but with much higher completeness across different target classes. After completing the SV phase, DESI began its main 5-year survey program in May 2021.
@@ 2.1 Targets
Details of the DESI target selection and different target classes are given in Myers et al. (2023); DESI Collaboration et al. (2024a, b); Cooper et al. (2023). Here, we provide a summary relevant to Milky Way science and necessary for the interpretation of the value-added catalogue.
@@ 2.2 High-level DESI target types
The DESI_TARGET bitmask encodes the overall DESI survey component for which each source was observed and also separates science targets from flux calibration standard stars and sky fibres. Table 3 lists the subset of DESI_TARGET bits corresponding to standard stars, MWS targets and secondary targets. These bits have the same meaning for SV1, SV2 and SV3, although the actual selections did change somewhat [^2] The magnitude range for STD_BRIGHT was changed from $1815\leq G<$ 18 in SV1 to $1816\leq G<$ 18 in SV2 and SV3. The colour criteria and the magnitude range of STD_FAINT were the same in SV1, SV2 and SV3.. The DESI_TARGET bitmask for a given SV stage is nonzero only for stars observed in that stage. For example, only stars observed in SV1 have a non-zero value of SV1_DESI_TARGET [^3] The VAC also has a small number (759) of rows with zero SV1_DESI_TARGET, SV2_DESI_TARGET and SV3_DESI_TARGET, but non-zero DESI_TARGET. These are fibres targeted as sky but having significant flux in them..
@@
STD_FAINT 33 Standard stars for dark/grey conditions
@@
GAIA_STD_FAINT 33 Gaia-based standard stars for dark/grey conditions
@@ 2.3 MWS survey targets
  • MWS_MAIN_BROAD class correspond to a magnitude limited selection in the range $1916<r<$ 19 without kinematic or colour cuts, supplemented by basic photometric quality requirements and was used in SV1 only  [^4] Several target classes (e.g. MWS_BROAD) have NORTH and SOUTH sub-categories (e.g. MWS_BROAD_NORTH and MWS_BROAD_SOUTH). Sources selected using Bock/Mosaic photometry in the northern section of the DESI survey footprint have the NORTH bit set, while those selected using DECam photometry in the southern section of the footprint have the SOUTH bit set. Since the MWS selection criteria were identical for the northern and southern parts of the footprint, these subclasses only indicate the origin of the input photometry..
@@
  • MWS_BHB comprises blue horizontal branch stars selected using the DESI Legacy Survey photometry and was used from SV1. We note that this category, which is described in (Cooper et al., 2023), is distinct from the BHB secondary target category defined for observations of more distant BHB candidates in the dark time (see below).
@@
For all the target classes listed above in Tables  6, 7, 8 we provide the number of objects in these classes observed in different surveys and programmes included in this VAC. Note that the same object may have more than one target class and may have been observed in multiple surveys.
In this paper we do not describe the selection function of the survey in detail, partially because this VAC is based on the early data release which is a combination of commissioning and science verification surveys and the target selection was changing throughout. However, for each survey and programme the information needed to reproduce the selection function or replay the target selection is available as part of the DESI EDR. Specifically:
@@ 2.4 Secondary targets
  • MWS_CALIB and BACKUP_CALIB are classes used during SV only to prioritise stars with data from one or more previous surveys, including APOGEE (Abdurro’uf et al., 2022), BOSS (Dawson et al., 2013), GALAH (De Silva et al., 2015) and Gaia-ESO (Gilmore et al., 2022). The MWS_CALIB stars have Gaia $G$-band magnitudes $1916<G<$ 19 while the BACKUP_CALIB stars have $G<16$.
@@ 2.5 Dedicated fields
Blank 98.0063 -4.8356  [^7] The field was designed to point at the Rosette Nebula, but was mistakenly centred at a negative declination of $\delta=-4.8$ instead of $\delta=4.8$. 1800.3 80718
@@ 4.1 RVS data products
We start by describing the individual RVS outputs of fitting results on coadded DESI spectra. The output of RVS (for a given survey, program and HEALPix pixel) consists of two files: an rvtab file that tabulates measurements by RVSpecFit and a rvmod file that contains the best-fit model for each spectrum.
@@
  • RVS_WARN is an integer bitmask set to zero if there are no warnings during the RVS fit. The first bit (or a bit number zero) of the bitmask is set if the CHISQ_TOT - CHISQ_C_TOT $<$ 50, which means that the stellar fit provides only a small improvement over the continuum-only fit. The second bit of RVS_WARN is set to one if the best-fit velocity is within 5 $\mathrm{\,km\,s^{-1}}$ of the radial velocity grid edges ($\pm$1500 $\mathrm{\,km\,s^{-1}}$); the third bit is set to one if the radial velocity error value is larger than 100 $\mathrm{\,km\,s^{-1}}$. The remaining bits of the bitmask are unused in this VAC.
@@ 5.3 Stellar parameter distribution
In Figure 5, we show the distributions of surface gravity and effective temperature (Kiel diagrams) for stars in the VAC, as measured by the RVS (left) and SP pipelines (right). The distributions are similar and show features that are expected, such as the red giant branch (extending from $\log\,g=4$, $T_{\mathrm{eff}}=5500$ K to $\log\,g=1$, $T_{\mathrm{eff}}=4000$ K), the blue horizontal branch ($\log\,g$ $\sim$3, $T_{\mathrm{eff}}>6,500$ K) and the main sequence ($\log\,g$ <4, $T_{\mathrm{eff}}$ <6,000 K). The shape of the main sequence below the turn-off, however, is not very realistic and is likely dominated by systematic effects from the stellar atmospheric grids used in our pipelines. Both pipelines show clustering around grid points in $\log\,g$ and $T_{\mathrm{eff}}$. In addition, the SP pipeline also shows problematic clustering at $T_{\mathrm{eff}}=5000$ K and $\log\,g=5$. Some of these SP artefacts mostly disappear when using stricter selections described in Appendix B.
@@
Figure 6: The distribution of iron abundances for stars in the MWS VAC as determined by the RVS and SP pipelines. The distribution shows the peak associated with thick disk stars at $\mathrm{\,[Fe/H]}$ =-0.5, the peak associated with GSE at $\mathrm{\,[Fe/H]}=-1.2$ and metal-poor stellar halo at lower metallicities. The distributions are shown with the bin size of 0.1 dex. Due to gridding issue associated with the RVS measurements, would the plot be made with much smaller bins (i.e. 0.01 dex), we would see peaks at location of template grid with the step of 0.25 dex.
@@ 5.6 Draco dwarf galaxy dedicated fields
By looking at the metallicity distribution of Draco member stars in DESI, we find a median metallicity of $\mathrm{\,[Fe/H]}\approx-2.25\pm 0.03$, which at face value seems to differ significantly (by $\sim 0.3$ dex) from the median metallicity of $-1.93\pm 0.01$ reported by Kirby et al. (2011). However, Kirby et al. (2011) measured a significant metallicity gradient in Draco of $\approx-0.73$ dex deg-1. The median angular distance of the member stars in the DESI sample from the centre of Draco is $\sim$ 0.25 degrees, as opposed to $\sim$ 0.1 degrees for the literature sample, owing to the large field of view of DESI. This metallicity gradient can, therefore, explain at least $\sim 0.1$ dex of the discrepancy between Kirby et al. (2011) and the DESI measurement. In fact, the median metallicity of the DESI members within 0.1 degrees is $-1.97$ dex perfectly consistent with previous measurements.
@@ Acknowledgments
All the data shown in the published graphs or necessary to reproduce them are available in a machine-readable form on Zenodo https://doi.org/10.5281/zenodo.10498911. The value-added catalogue itself is available at https://data.desi.lbl.gov/doc/releases/edr/vac/mws/.
@@ Appendix A Selection criteria of the main MWS targets in SV3
In this appendix we provide the selection criteria for the SV3 targets. The selection code is also available at https://desitarget.readthedocs.io/en/latest/_modules/desitarget/sv3/sv3_cuts.html. We remark that the EDR covers DESI observations prior to the main survey. These targets were selected using Gaia DR2 photometry and astrometry, while the main survey selection uses Gaia DR3 astrometry. These selection criteria are slightly different from the main survey selection criteria described in Cooper et al. (2023).
@@
SUCCESS Boolean flag indicating whether the code has likely produced useful results
@@
HEALPIX HEALPix pixel containing this location at NSIDE=64 in the NESTED scheme